# Reweighted Spectral Partitioning Works: Bounds for Special Graph Classes

### v2

@@ 1 Introduction
Partitioning a graph into smaller pieces is a fundamental type of problem that arises in a variety of fields. Applications of graph partitioning include divide and conquer algorithms, dynamic programming algorithms [RH01], graph drawing [BPP22], FPGA compilation [CCW21], image processing, VLSI design, compute parallelization, bioinformatics, simulation, road networks, social networks [BMS+16], graph processing [OFS17], and even air traffic control methods [FGBJ17].
@@
A subset of vertices $S\subset V$ is said to be an $\alpha$-vertex-separator if there are disjoint sets $A\cup B=V\setminus S$ such that $|A|,|B|\leq\alpha\cdot n$. Similarly, a subset of edges $S\subset E$ is an $\alpha$-edge-separator if there are disjoint sets $A\cup B=V$ such that $|A|,|B|\leq\alpha\cdot n$. In both cases, the important quantity is usually the size $|S|$ of the separator. If $\alpha$ is omitted, it is assumed to be $\frac{2}{3}$. This is usually a reasonable assumption, since if a every induced subgraph $H$ of $G$ has an $\alpha$-vertex-separator of size at most $f(|V(H)|)$ (for some non-decreasing function $f$), then $G$ has a $2/3$-vertex-separator of size at most $\lceil\log_{\alpha}(2/3)\rceil f(|V(G)|)$. For simplicity, we will use the term separator to refer to a $\frac{2}{3}$-vertex-separator. Finding the minimum-size separator of a graph is NP-hard even for graphs of maximum degree $3$, and even up to a non-trivial additive approximation [BJ92]. However, small separators are known to exist for a large number of common graph classes. Most notably, the planar separator theorem states that a planar graph with $n$ vertices has a separator consisting of $O(\sqrt{n})$ vertices [Ung51, LT79]. We list this theorem, along with many other results for separators on various graph classes, in Table 1. See Appendix A for definitions of each class.
@@
(5,3pt,3pt)[5pt]—l—c—c—c——c—c—c—c—c—c—c— Graph class & Separator Size Time Refs.
@@
Genus-$g$ $\operatorname{\mathcal{O}}(\sqrt{g\,n})$ $\operatorname{\mathcal{O}}(n)$ [^1] Assumes an embedding of $G$ on a genus-$g$ surface is given. Computing such an embedding (or even the genus) is NP-complete [Tho89]. [GHT84]
@@
Table 1: Best-known sizes of $2/3$-vertex-separators for several graph classes, as well as the time required to compute the separator. Many of these results require additional information besides the graph to compute the separator in the stated time complexity (see the footnotes). $n$ is always the number of vertices, and $m$ is always the number of edges.
@@ 1.1.1 Spectral Partitioning
(5,3pt,3pt)[5pt]—l—c—c——c—c—c—c—c—c—c—c— Graph class & $\lambda_{2}(G)\lesssim$ Refs.
@@
Genus-$g$ $\dfrac{\Delta\,g(\log g)^{2}}{n}$ [BLR10, LS10]
Triangulated [^1] Due to a subtle issue, this bound does not apply to general genus-$g$ graphs, which is what the references claim. We will amend this issue as part of our work. genus-$g$ $\dfrac{\text{poly}(\Delta)\cdot g}{n}$ [Kel04, Kel06b, Kel06a]
@@ 1.1.2 Reweighted Spectral Partitioning
Boyd, Diaconis, and Xiao also realize their optimization problem as a Semidefinite Program (therefore allowing it to be approximately solved in polynomial time via e.g. [ZL18]), and in the process they derive a dual formulation (better stated and properly extracted in follow-up work by Roch [Roc05]):
@@
If $n$ is even, assign $y$ to be $\frac{2}{n}$ for every vertex, and assign $f$ to be $-1$ for half the vertices and $1$ for the other half of the vertices, so then $\gamma^{(1)}(G)\leq 2\lesssim 1$. If $n$ is odd (and at least $3$), assign $y$ to be $\frac{2}{n-1}$ for every vertex, and assign $f$ to be an even split of $-1$ and $1$ for all but one vertex, which is assigned 00, resulting in $\gamma^{(1)}(G)\leq 2+\frac{2}{n-1}\lesssim 1$. ∎
@@ 1.2 Results and Organization
We will give bounds on $\lambda_{2}^{*}$, $\gamma^{(1)}$, and $\gamma^{(d)}$ for a large variety of graph classes, summarized in Table 3 and Table 4. In turn, these imply bounds on (vertex) separator sizes, as well as bounds on the resulting separators from reweighted spectral partitioning, summarized in Table 5. Rather surprisingly, this implies some improved (non-constructive) separator bounds for some geometric graph classes (likely since typical separator bounds for geometric graph classes do not focus on $\frac{2}{3}$-separators directly). Most of the bounds on $\lambda_{2}^{*}$ will primarily adapt existing machinery of bounds for $\lambda_{2}$. However, for genus-$g$ graphs, we are able to make some notable improvements to the techniques, and our bound on $\lambda_{2}^{*}$ also implies a new bound on $\lambda_{2}$. These techniques are the topic of Section 3.2.
@@
(2,1pt,1pt)[4pt]—c—c——c—c—c—c—c— Graph class & $\frac{\lambda_{2}}{\Delta}\leq\lambda_{2}^{*}\leq\gamma^{(1)}\lesssim$
@@
(2,1pt,1pt)[4pt]—c—c——c—c—c— Graph class & $\frac{\lambda_{2}}{\Delta}\leq\lambda_{2}^{*}\leq\gamma^{(d)}\lesssim$
@@
(3,1pt,1pt)[4pt]—c—c—c——c—c—c—c—c—c—c— Graph class & Separator size $\lesssim$ Poly-time separator size $\lesssim$
@@
Genus-$g$ $\min\{\log g,\sqrt{\log\Delta}\}\sqrt{gn}$ $\min\{(\log g)^{2},\log\Delta\}\sqrt{gn}$
@@ 2 Detailed Background
In this section, we give some more detailed background that is important to our results. We will first discuss the history of bounds on $\lambda_{2}(G)$ for a few graph classes in more detail. Then we will discuss two groups of tools used for some of those bounds.
@@ 2.1 Bounds on $\lambda_{2}(G)$ for Special Graph Classes
For a triangulated genus-$g$ graph $G$ with $n$ vertices and maximum degree $\Delta$, the second smallest eigenvalue of the Laplacian matrix of $G$ has $\lambda_{2}(G)\lesssim\frac{\text{poly}(\Delta)\cdot g}{n}$, where $\text{poly}(\Delta)$ denotes a polynomial in $\Delta$.
This result is of particular interest since the proof is entirely geometric, and it was essentially the last result of this kind to have a geometric proof. As mentioned earlier, the references for this result claim that it applies to general genus-$g$ graphs. This is due to the fact that any genus-$g$ graph can be triangulated by adding edges (i.e., without adding new vertices), and adding edges only increases $\lambda_{2}$. The issue with this logic is that adding edges naively could also very significantly increase $\Delta$, even to the order of $n$. A method of triangulation by adding edges that asymptotically preserves $\Delta$ is known for planar graphs [KB92, Kan93, KB97], but, to the best of our knowledge, no such result is known for genus-$g$ graphs. We will later bypass this issue with a careful triangulation that also adds vertices.
@@ 2.2 Extremal Spread
Note that, by definition, $\overline{s}_{2}(G)\geq\overline{s}_{1}(G)$.
@@
For a genus-$g$ graph $G$ with $n$ vertices, if $n\geq 3\sqrt{g}$, then
@@ 2.3 Non-Expansive Metric Embeddings
Let $(X,d)$ be a metric space. The (weak) modulus of padded decomposability of $(X,d)$ is defined as the smallest value of $\alpha$ so that for every $\Delta>0$, there exists a distribution $\mu$ of partitions of $X$ with parts of diameter at most $\Delta$, and
@@
For $p\geq 1$, there exists a (universal) constant $C_{p}\geq 1$ (depending only on $p$), so that for every metric space $(X,d)$, there exists a non-expansive metric embedding $f:X\to\mathbb{R}$ with average $p$-distortion at most $C_{p}\left[\alpha(X,d)\right]^{p}$.
@@
Let $G$ be a $K_{h}$-minor-free graph. Then $\alpha(G)\in\operatorname{\mathcal{O}}(\log h)$. Consequently, for any set of edge or vertex weights $\omega$ and an induced shortest-path distance metric $d_{\omega}$, there exists a non-expansive metric embedding function $f:(V(G),d_{\omega})\to\mathbb{R}$ with average $2$-distortion at most $\operatorname{\mathcal{O}}((\log h)^{2})$.
@@ 3 Bounds via Geometric Intersection Graphs
Let $B_{1},\dots,B_{n}$ be a collection of balls in $\mathbb{R}^{d}$ with centres $c_{1},\dots,c_{n}$, so that no point $x\in\mathbb{R}^{d}$ is contained in $\left\lceil\frac{n}{2}\right\rceil$ of the balls. Then there is an homeomorphism $\alpha$ from $\mathbb{R}^{d}$ to a subset of the sphere $S^{d}$,  [^1] We use the standard topological convention of $S^{d}$ as the $d$-dimensional surface of the $(d+1)$-dimensional unit ball embeddable in $\mathbb{R}^{d+1}$. This differs from the notation of Spielman and Teng, who used $S^{d}$ to denote the $(d-1)$-dimensional surface of the $d$-dimensional unit ball. so that $\alpha(B_{i})$ is exactly a geodesic ball in $S^{d}$ with center $\alpha(c_{i})$, and moreover so that the centroid of the values $\alpha(c_{1}),\dots,\alpha(c_{n})$, in the natural representation of $S^{d}$ as surface of the unit ball of $d+1$ dimensions, is exactly the origin.
@@
By Theorem 3.3 and Theorem 3.4, there exists a representation of $G$ as a circle packing on the sphere $S^{2}\subset\mathbb{R}^{3}$. Let the centers be given by $f:V\to S^{2}\subset\mathbb{R}^{3}$ (i.e., $||f(v)||_{2}^{2}=1$ for each $v\in V$) and let the radii be given by $s:V\to\mathbb{R}_{\geq 0}$. We claim first that $f,s$ form a feasible solution to $\dot{\gamma}^{(3)}(G)$: The constraint $\sum_{v\in V}f(v)=\overline{0}$ is exactly the statement that the centroid of the centres is the origin, which is given by Theorem 3.4. The other constraints follow from the facts that two balls intersect if and only if they share an edge, and that $||f(u)-f(v)||_{2}$ is bounded above by the geodesic distance along $S^{2}$ between $f(u)$ and $f(v)$. Next, we bound the objective value: It follows from the statement of Theorem 3.3 that the ply of all points in $S^{2}$ except for a set of measure 00 is at most $1$. Hence, the total area of all balls is bounded by the area of $S^{2}$ itself, which is known to be $4\pi$. The area of the ball on $S^{2}$ for vertex $v$ is bounded above by $\pi(s(v))^{2}$, since $s(v)$ is the (geodesic) radius of the ball, so $\sum_{v\in V}\pi(s(v))^{2}\leq 4\pi$. Therefore, we get a bound on the objective value of
@@ 3.1 Geometric Intersection Graphs and Nearest-Neighbour Graphs
Let $A_{d}$ denote the area of the $d$-sphere, and let $V_{d}$ denote the volume of the $d$-ball (both of these are $d$-dimensional measures). [^2] We again use the topological convention of the $d$-sphere $S^{d}$ as the $d$-dimensional surface of the $(d+1)$-dimensional unit ball. Let $G$ be a $d$-dimensional $k$-ply neighbourhood system with $n\geq d+1$ vertices and maximum degree $\Delta$. Then,
@@
The proof is very similar to that of Theorem 3.5, and again is based on the methods of Spielman and Teng [ST96, ST07, proof of Theorem 5.1]. If we can show that $\dot{\gamma}^{(d+1)}(G)\leq\left(\frac{A_{d}}{V_{d}}\cdot\frac{k}{n}\right)^{\frac{2}{d}}$, then the other inequalities follow from previously-stated theorems, or from Gautschi’s inequality [Gau59] applied to $\frac{A_{d}}{V_{d}}$ (via the closed-form expressions of $A_{d}$ and $V_{d}$ using the Gamma function). By Theorem 3.4, we may realize $G$ as the geometric intersection graph of a set of geodesic balls on $S^{d}$, so that every point in $S^{d}$ is contained in at most $k$ of these geodesic balls almost surely, and so that the centroid of the centres of the geodesic balls in $\mathbb{R}^{d+1}$ is exactly the origin. Denote the geodesic balls as $B_{1}^{\prime},\dots,B_{n}^{\prime}$ with (geodesic) radii $r_{1}^{\prime},\dots,r_{n}^{\prime}$ and centres $c_{1}^{\prime},\dots,c_{n}^{\prime}\in S^{d}$, and let $\text{volume}(B_{i}^{\prime})$ denote the ($d$-dimensional) volume of the geodesic ball $B_{i}^{\prime}$ on $S^{d}$. Let $r_{i}^{\prime\prime}$ denote the maximum euclidean distance from $c_{i}^{\prime}$ to some other point in $B_{i}^{\prime}$ using the natural embedding of the unit $d$-ball into $\mathbb{R}^{d+1}$-dimensional space as the surface of the $(d+1)$-dimensional unit ball. We obtain $r_{i}^{\prime\prime}\leq r_{i}^{\prime}$. See Figure 2 for a low-dimensional example of these quantities on a geodesic ball. Then,
@@
Moreover, by the definition of $\{r_{i}^{\prime\prime}\}$, if $B_{i}^{\prime}$ and $B_{j}^{\prime}$ intersect (i.e. there is an edge $ij$ in $G$), then $r_{i}^{\prime\prime}+r_{j}^{\prime\prime}\geq||c_{i}-c_{j}||_{2}$. Thus, we have satisfied the constraints of $\dot{\gamma}^{(d)}(G)$. Note that $||c_{i}||_{2}^{2}=1$ when the points are represented as the surface of the $(d+1)$-dimensional unit ball. Hence,
@@
Let $G$ be a $d$-dimensional $k$-ply neighbourhood system with maximum degree $\Delta$, provided as a set of $d$-dimensional coordinates and radii for each vertex. Then, in polynomial time, we can compute a balanced $2/3$-vertex-separator of size $\operatorname{\mathcal{O}}\left(\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}\right)$.
@@
Let $G$ be a $d$-dimensional $k$-nearest neighbour graph with maximum degree $\Delta$, provided as a set of $d$-dimensional coordinates and a value $k$. Then, in polynomial time, we can compute a balanced $2/3$-vertex-separator of size $\operatorname{\mathcal{O}}\left(\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}\right)$.

### v3

@@ 1 Introduction
Partitioning a graph into smaller pieces is a fundamental type of problem that arises in a variety of fields. Applications of graph partitioning include divide and conquer algorithms, dynamic programming algorithms [RH01], graph drawing [BPP22], geometrically embedding graphs [SSX25], FPGA compilation [CCW21], image processing, VLSI design, compute parallelization, bioinformatics, simulation, road networks, social networks [BMS+16], graph processing [OFS17], and even air traffic control methods [FGBJ17].
@@
[…]
@@
(5,3pt,3pt)[5pt]—l—c—c—c——c—c—c—c—c—c—c— Graph class Separator Size Time Refs.
@@
[…]
@@
Table 1: Best-known sizes of 2/3 2/32 / 3-vertex-separators for several graph classes, as well as the time required to compute the separator. Many of these results require additional information besides the graph to compute the separator in the stated time complexity (see the footnotes). $n$ is always the number of vertices, and $m$ is always the number of edges.
@@ 1.1.1 Spectral Partitioning
(5,3pt,3pt)[5pt]—l—c—c——c—c—c—c—c—c—c—c— Graph class $\lambda_{2}(G)\lesssim$ Refs.
@@
[…]
Triangulated [^1] Due to a subtle issue, this bound does not apply to general $g$ graphs, which is what the references claim. We will amend this issue as part of our work. $g$ $\dfrac{\text{poly}(\Delta)\cdot g}{n}$ [Kel04, Kel06b, Kel06a]
@@ 1.1.2 Reweighted Spectral Partitioning
Boyd, Diaconis, and Xiao also realize their optimization problem as a Semidefinite Program meeting Slater’s condition (therefore allowing it to be approximately solved in polynomial time via e.g. [ZL18]), and in the process they derive a dual formulation (better stated and properly extracted in follow-up work by Roch [Roc05]):
@@
If $n$ is even, assign $y$ to be $\frac{2}{n}$ for every vertex, and assign $f$ to be 1-1- 1 for half the vertices and 111 for the other half of the vertices, so then $\gamma^{(1)}(G)\leq 2\lesssim 1$. If $n$ is odd (and at least 333), assign $y$ to be $\frac{2}{n-1}$ for every vertex, and assign $f$ to be an even split of 1-1- 1 and 111 for all but one vertex, which is assigned 0, resulting in $\gamma^{(1)}(G)\leq 2+\frac{2}{n-1}\lesssim 1$. ∎
@@ 1.2 Results and Organization
We will give bounds on $\lambda_{2}^{*}$, $\gamma^{(1)}$, and $\gamma^{(d)}$ for a large variety of graph classes, summarized in Table 3 and Table 4. In turn, these imply bounds on (vertex) separator sizes, as well as bounds on the resulting separators from reweighted spectral partitioning, summarized in Table 5. This also slightly improves the dependence on dimension for separator bounds in some geometric graph classes. Most of our bounds on $\lambda_{2}^{*}$ will primarily adapt existing machinery of bounds for $\lambda_{2}$. However, for $g$ graphs, we are able to make some notable improvements to the techniques, and our bound on $\lambda_{2}^{*}$ also implies a new bound on $\lambda_{2}$. These techniques are the topic of Section 3.2. The main new tool is the use of something we call “uniform” shallow minors, although these are used in a very different manner compared to previous work leveraging forbidden shallow minors.
@@
(2,1pt,1pt)[4pt]—c—c——c—c—c—c—c— Graph class $\frac{\lambda_{2}}{\Delta}\leq\lambda_{2}^{*}\leq\gamma^{(1)}\lesssim$
@@
(2,1pt,1pt)[4pt]—c—c——c—c—c— Graph class $\frac{\lambda_{2}}{\Delta}\leq\lambda_{2}^{*}\leq\gamma^{(d)}\lesssim$
@@
(3,1pt,1pt)[4pt]—c—c—c——c—c—c—c—c—c—c— Graph class Separator size $\lesssim$ Poly-time separator size $\lesssim$
@@
[…]
@@ 2 Detailed Background
In this section, we give some more detailed background that is important to our methods. We will first discuss the history of bounds on $\lambda_{2}(G)$ for a few graph classes in more detail. Then we will discuss two groups of tools used for some of those bounds.
@@ 2.1 Bounds on $\lambda_{2}(G)$ for Special Graph Classes
For a triangulated $g$ graph $G$ with $n$ vertices and maximum degree $\Delta$, the second smallest eigenvalue of the Laplacian matrix of $G$ has $\lambda_{2}(G)\lesssim\frac{\text{poly}(\Delta)\cdot g}{n}$, where $\text{poly}(\Delta)$ denotes a polynomial in $\Delta$.
This result is of particular interest since the proof is entirely geometric, and it was essentially the last result of this kind to have a geometric proof. As mentioned earlier, the references for this result claim that it applies to general $g$ graphs. This is due to the fact that any $g$ graph can be triangulated by adding edges (i.e., without adding new vertices), and adding edges only increases $\lambda_{2}$. The issue with this logic is that adding edges naively could also very significantly increase $\Delta$, even to the order of $n$. A method of triangulation by adding edges that asymptotically preserves $\Delta$ is known for planar graphs [KB92, Kan93, KB97], but, to the best of our knowledge, no such result is known for $g$ graphs. We will later bypass this issue with a careful triangulation that also adds vertices.
@@ 2.2 Extremal Spread
{TAB}
(2,1pt,1pt)[4pt]—c—c—c——c—c—c—c—c— Graph class $\overline{s}_{2}(G)\gtrsim$ Refs.
[…]
$K_{h}$-minor-free $n\geq 4h$ $\dfrac{n^{2}}{h^{3/2}}$ [BLR10, Theorems 3.9 and 3.11]
$K_{h}$-minor-free $n\gtrsim h\sqrt{\log h}$ $\dfrac{n^{2}}{h\sqrt{\log h}}$ [BLR10, Theorems 3.9 and 3.11]
RIG over $K_{h}$-minor-free $\dfrac{n^{2}}{h\sqrt{\Delta\log h}}$ [Lee17, Corollary 5.4]
Table 6: Asymptotic lower bounds on the $L^{2}$-extremal spread $\overline{s}_{2}(G)$ for various graph classes. $n$ is always the number of vertices, and $\Delta$ is the maximum degree.
Note that, by definition, $\overline{s}_{2}(G)\geq\overline{s}_{1}(G)$. A number of bounds for $\overline{s}_{2}$ are known, which we will review. A summary can be found in Table 6.
@@
For a $g$ graph $G$ with $n$ vertices, if $n\geq 3\sqrt{g}$, then
@@ 2.3 Non-Expansive Metric Embeddings
[…]
@@
[…]
@@
Let $G$ be a $K_{h}$-minor-free graph. Then $\alpha(G)\in\operatorname{\mathcal{O}}(\log h)$. Consequently, for any set of edge or vertex weights $\omega$ and an induced shortest-path distance metric $d_{\omega}$, there exists a non-expansive metric embedding function $f:(V(G),d_{\omega})\to\mathbb{R}$ with average 222-distortion at most $\operatorname{\mathcal{O}}((\log h)^{2})$.
@@ 3 Bounds via Geometric Intersection Graphs
[…]
@@
By Theorem 3.3 and Theorem 3.4, there exists a representation of $G$ as a circle packing on the sphere $S^{2}\subset\mathbb{R}^{3}$. Let the centers be given by $f:V\to S^{2}\subset\mathbb{R}^{3}$ (i.e., $||f(v)||_{2}^{2}=1$ for each $v\in V$) and let the radii be given by $s:V\to\mathbb{R}_{\geq 0}$. We claim first that $f,s$ form a feasible solution to $\dot{\gamma}^{(3)}(G)$: The constraint $\sum_{v\in V}f(v)=\overline{0}$ is exactly the statement that the centroid of the centres is the origin, which is given by Theorem 3.4. The other constraints follow from the facts that two balls intersect if and only if they share an edge, and that $||f(u)-f(v)||_{2}$ is bounded above by the geodesic distance along $S^{2}$ between $f(u)$ and $f(v)$. Next, we bound the objective value: It follows from the statement of Theorem 3.3 that the ply of all points in $S^{2}$ except for a set of measure 0 is at most 111. Hence, the total area of all balls is bounded by the area of $S^{2}$ itself, which is known to be $4\pi$. The area of the ball on $S^{2}$ for vertex $v$ is bounded above by $\pi(s(v))^{2}$, since $s(v)$ is the (geodesic) radius of the ball, so $\sum_{v\in V}\pi(s(v))^{2}\leq 4\pi$. Therefore, we get a bound on the objective value of
@@ 3.1 Geometric Intersection Graphs and Nearest-Neighbour Graphs
[…]
@@
[…]
@@
[…]
@@
Let $G$ be a $d$-dimensional $k$-ply neighbourhood system with maximum degree $\Delta$, provided as a set of $d$-dimensional coordinates and radii for each vertex. Then, in polynomial time, we can compute a balanced 2/32/32 / 3-vertex-separator of size $\operatorname{\mathcal{O}}\left(\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}\right)$.
@@
Let $G$ be a $d$-dimensional $k$-nearest neighbour graph with maximum degree $\Delta$, provided as a set of $d$-dimensional coordinates and a value $k$. Then, in polynomial time, we can compute a balanced 2/32/32 / 3-vertex-separator of size $\operatorname{\mathcal{O}}\left(\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}\right)$.